OpenStudy (anonymous):

Suppose a bag of marbles has 4 green, 2 red, 5 yellow, 1 brown, and 7 blue marbles. What is the probability of picking a yellow marble, replacing it, and picking another yellow marble?

OpenStudy (amistre64):

how many yellows we got?

OpenStudy (anonymous):

5

OpenStudy (amistre64):

and how many marbles we got altogether?

OpenStudy (amistre64):

looks like 19 i think

OpenStudy (amistre64):

so P(y) = 5/19 by replacing the marble, none of the values change in the next pick soo P(y and y) = P(y) * P(y) = (5/19)^2